### v1

# How many simulations do we need for simulation-based inference in cosmology?

Abstract
How many simulations do we need to train machine learning methods to extract information available from summary statistics of the cosmological density field? Neural methods have shown the potential to extract non-linear information available from cosmological data. Success depends critically on having sufficient simulations for training the networks and appropriate network architectures. In the first detailed convergence study of neural network training for cosmological inference, we show that currently available simulation suites, such as the Quijote Latin Hypercube(LH) with 2000 simulations, do not provide sufficient training data for a generic neural network to reach the optimal regime, even for the dark matter power spectrum, and in an idealized case. We discover an empirical neural scaling law that predicts how much information a neural network can extract from a highly informative summary statistic, the dark matter power spectrum, as a function of the number of simulations used to train the network, for a wide range of architectures and hyperparameters. We combine this result with the Cramer-Rao information bound to forecast the number of training simulations needed for near-optimal information extraction. To verify our method we created the largest publicly released simulation data set in cosmology, the Big Sobol Sequence(BSQ), consisting of 32,768 $\Lambda$CDM n-body simulations uniformly covering the $\Lambda$CDM parameter space. Our method enables efficient planning of simulation campaigns for machine learning applications in cosmology, while the BSQ dataset provides an unprecedented resource for studying the convergence behavior of neural networks in cosmological parameter inference. Our results suggest that new large simulation suites or new training approaches will be necessary to achieve information-optimal parameter inference from non-linear simulations.
preprint: APS/123-QED
I Introduction
The standard model of cosmology[1, 2] has emerged from compelling observational evidence across multiple scales and probes. Dark matter, first indicated by galaxy rotation curves[3, 4, 5] and later confirmed through weak gravitational lensing[6], comprises approximately 25% of the universe’s energy content, as measured through cosmic microwave background (CMB) observations from WMAP and Planck[7, 8]. CMB data reveals that ordinary matter and dark matter together account for only about 30% of the critical density needed for a flat universe. This finding, combined with the discovery of cosmic acceleration[9, 10], points to the existence of dark energy—represented in $\Lambda$CDM by a cosmological constant ($\Lambda$)—which makes up the remaining 70% of the cosmic energy budget.
Thus, a major effort in modern cosmology is directed towards inferring the cosmological parameters from astronomical observations and testing the fundamental tenets of $\Lambda$CDM. This endeavor is supported by increasingly precise numerical simulations[11] and forthcoming data from several major cosmological surveys, including Euclid[12], DESI[13], Vera Rubin Observatory[14], the Simons Observatory[15], and CMB-S4[16].
While sophisticated methods for cosmological parameter inference have dramatically reduced our uncertainty about cosmological parameters in recent decades[17, 18, 19, 20, 21, 22], current and upcoming surveys present new methodological challenges. The traditional approach compares summary statistics—such as n-point correlation functions, number counts, and wavelet statistics—against theoretical predictions or numerical simulations[23, 24, 25, 26]. These methods typically assume an explicit form of the likelihood function, often Gaussian, but this assumption becomes problematic in the non-linear regime where the true likelihood form may not be known, especially in the regime of a few modes and complex modeling. For example, even when the number of modes is large, non-linear evolution can lead to strong correlations, slowing the asymptotic approach to the Gaussian form. Even when good approximations do exist, standard Bayesian inference techniques like Markov Chain Monte Carlo (MCMC) can face convergence challenges[27].
Implicit Inference (II), or Simulation-Based Inference (SBI), has emerged as a promising alternative that learns the joint statistics of data and parameters directly from simulated pairs, avoiding the need for explicit likelihood assumptions. SBI has demonstrated both accuracy and computational efficiency across various applications[28, 29, 30, 31, 32, 33, 34]. Neural networks play two crucial roles in modern SBI approaches: generating or compressing informative summary statistics and representing statistical distributions of likelihood or posterior. Information Maximizing Neural Networks (IMNN) exemplify the first role, focusing on Fisher information maximization[35], while (ensembles of) normalizing flows[36, 37] and diffusion models[38, 39, 40] excel at the second.
Given the centrality of trained neural networks to this powerful approach, a key question emerges: How many simulations are needed to achieve optimal results? While one could generate simulations until the extracted information plateaus, this brute-force approach is computationally expensive or even intractable. Our work presents the first comprehensive analysis of neural network convergence in cosmological information extraction and provides a predictive framework to determine the required number of simulations.
For our investigation, we focus on the power spectrum $P(k)$ as our summary statistic. This choice is motivated by the fundamental role of $P(k)$ in cosmology: at large scales and high redshifts, where the universe remains approximately Gaussian and homogeneous, $P(k)$ captures the complete statistical description[41, 27, 42, 43, 24]. While current three-dimensional surveys access significant non-Gaussian information at smaller scales and lower redshifts, any method claiming to extract cosmological information should, at minimum, demonstrate proficiency with $P(k)$ in the non-linear regime. Our findings reveal that even for this basic case, the required training set size is substantial, suggesting that simulation requirements may limit the amount of information extracted from cosmological data and will be a significant challenge for brute-force neural network-based analyses.
The remainder of this paper is organized as follows. Section II describes our simulation suite. Section III presents methods for measuring parameter information content in cosmological datasets. Section IV details our neural network methodology. In section V we present our method for estimating the number of required training sets to asymptote to the Cramer Rao information bound. In order to test these predictions we present the Big Sobol Sequence (BSQ) in section V.1 comprising 32,768 simulations designed specifically for testing neural network convergence and quantifying information content. Our analysis includes careful consideration of the important topic of network architecture selection (VIII.1) and hyperparameter optimization (VIII.1). We discuss our findings and conclude in section VI.
II The Quijote Simulation Suite
The Quijote Simulation Suite is a collection of more than 88,000 full N-body dark matter simulations designed to probe the large-scale structure of the universe[11]. Using the TreePM code GADGET-III which is an improved version of GADGET-II[44], these simulations span more than 40,000 cosmological models, exploring a parameter space defined by ${\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8},M_{\nu},w}$. Each simulation models the evolution of dark matter in a cubic volume of 1$(Gpc/h)^{3}$, computed at three spatial resolutions: $256^{3}$, $512^{3}$, and $1024^{3}$ grid points.
The suite includes several complementary subsets designed for different analysis approaches. For instance, the Latin Hypercube (LH) subset consists of 2,000 simulations sampling five parameters ${\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}}$, specifically configured for machine learning applications. These parameters are sampled uniformly within ranges centered around the Planck 2018 best-fit values[2]. Additionally, Quijote provides 15,000 fiducial simulations and 500 pairs of finite difference simulations for robust statistical analysis.For our investigation of simulation-based inference using neural networks, we focus on the matter power spectrum $P(k)$ measured at redshift $z=0$. We analyze scales up to $k_{max}=0.804,h/\text{Mpc}$, corresponding to the $0.25$ times Nyquist frequency of a $1024^{3}$ grid in our 1$(Gpc/h)^{3}$ volume. The power spectrum is evaluated at multiples of the fundamental frequency, yielding a 128-dimensional measurement vector over this $k$ range.
III Information Content: Fisher Information and Forecasts
The Fisher information formalism [45, 28] provides a framework for quantifying the information about cosmological parameters $\theta$ contained in data $d(\theta)$. This approach yields insights into maximum likelihood parameter estimates and their uncertainties. For unbiased estimators, the Cramér-Rao information inequality establishes a fundamental lower bound on parameter variance
$\text{Var}(\theta)\geq(F^{-1})_{\theta\theta},$ (1)
where $F$ is the Fisher information matrix, defined as: $F=-\langle\nabla_{\theta}\nabla^{T}_{\theta}\mathcal{L}\rangle=\langle\nabla_{\theta}\mathcal{L}\nabla^{T}_{\theta}\mathcal{L}\rangle$, and $\mathcal{L}$ is the log-likelihood of the statistic. Here, the expectation value is taken over different realizations at fixed $\theta$, and the Fisher matrix in equation (1) is evaluated at the fiducial parameter value $\theta=\theta_{\star}$.
In the asymptotic limit of a large number of modes, the power spectrum likelihood is approximately Gaussian, so the log-likelihood takes the form
$\mathcal{L}=-\frac{1}{2}(d-\mu(\theta))^{T}C^{-1}(d-\mu(\theta))-\frac{1}{2}\ln|2\pi C|,$ (2)
leading to a Fisher information matrix[46] of
$F=\nabla_{\theta}\mu^{T}C^{-1}\nabla_{\theta}\mu.$ (3)
The convergence of both $\nabla_{\theta}\mu$ and the covariance matrix $C$ in equation (3) depends critically on the number of simulations. The Quijote suite provides sufficient simulations for convergence, with 500 pairs of finite difference simulations and 15,000 fiducial simulations.
At the maximum likelihood point
$\nabla_{\theta}\mathcal{L}|_{\textbf{ML}}=\nabla{\theta}\mu^{T}|_{\textbf{ML}}C^{-1}(d-\mu(\theta_{\textbf{ML}}))=0.$
A Taylor expansion of this score function around the fiducial point $\theta_{\star}$ yields the Quasi-Maximum Likelihood estimator
$\theta_{\textbf{ML}}=\theta_{\star}+F^{-1}\nabla_{\theta}\mu^{T}({\theta_{\star}})C^{-1}(d-\mu(\theta_{\star}))$ (4)
at leading order in $\theta_{\textbf{ML}}-\theta_{\star}$. For parameters far from the fiducial point, equation (4) must be iterated until convergence, requiring covariance matrix evaluation at each intermediate point. However, this iteration is computationally prohibitive due to the large number of simulations required at each step.
In our analysis, we use the power spectrum $P(k)$ from the Quijote suite to establish optimal information bounds. The Fisher information matrix $F_{P(k)}$ is computed using the suite’s 15,000 fiducial simulations for the covariance matrix $C$ and 500 pairs of finite difference simulations for the derivative terms $\nabla_{\theta}\mu$. This framework provides a benchmark against which we can evaluate the performance of neural network-based inference methods, which we discuss in the following section.
IV Neural Estimator
Neural networks have become increasingly important tools in astrophysics and cosmology for extracting underlying parameters from simulated or observational data, owing to their ability to learn complex inverse models through chaining linear operations and nonlinear activation functions, combined with their differentiability, enabling efficient training with stochastic gradient descent. Developing effective models requires advances in multiple areas: architecture design, hyperparameter optimization, weight and bias initialization, and the relationship between network structure and the underlying physics of the data. In the following, we explore the application of neural networks in inferring cosmological parameters from power spectra $P(k)$ with particular attention to architecture selection, training, and hyperparameter tuning.
IV.1 Selection of the neural architecture
Multi-layer Perceptrons (MLPs) demonstrate superior performance compared to linear regression models for nonlinear parameter estimation tasks. As fully connected feed-forward neural networks with multiple hidden layers and nonlinear activation functions, MLPs excel at capturing complex relationships in data. We employ an MLP to infer cosmological parameters from simulated power spectra $P(k)$. After extensive architecture exploration, we developed a model (Figure 1) consisting of 6 hidden layers, each followed by LeakyReLU(0.5) activation. The hidden layer dimensions follow an expanding-contracting pattern $(128\rightarrow 512\rightarrow 1024\rightarrow 1024\rightarrow 512\rightarrow 128)$, which improves parameter estimation performance. We will study the effect of variations in the architecture in subsection VIII.1.
For neural network training, we divide the 2000 Latin Hypercube simulations into training (1500), validation (300), and test (200) sets. To manage memory constraints during training, we process the data in batches of 16. All power spectra are normalized before being input into the model to improve convergence. We use the Adam optimizer with an initial learning rate of 0.001 and momentum ($\beta$) of 0.9 to minimize the log MSE loss $\mathcal{L}$ between the true parameter $\theta$ and inferred parameter $\hat{\theta}$
$\mathcal{L}$ $=\sum_{i}\ln(MSE)_{i}$
$=\sum_{i}\ln\left(\frac{1}{N_{batch}}\sum_{n=0}^{N_{batch}}\big{[}\theta_{i}^{(n)}-\hat{\theta}_{i}^{(n)}\big{]}^{2}\right)$ (5)
where $i$ denotes different parameter classes. The learning rate is reduced by a factor $\gamma=0.9$ every 10 epochs using PyTorch’s StepLR scheduler. These hyperparameters have been chosen after an extensive hyperparameter search, as shown in VIII.1. Training the model on the 2000 Latin Hypercube $P(k)$ samples for 500 epochs required approximately one hour on a V100 GPU, with model weights at the lowest validation loss retained for inference.
Figure 1: An illustration of PowerSpectraNet architecture. This infers cosmological parameters $\hat{\theta}:\{\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}\}$ from the Power Spectrum $P(k)$. The dimension of each layer is mentioned at the bottom of the corresponding layer.
V Convergence Analysis and Scaling to estimate simulation requirements
To understand how many simulations are required for optimal parameter inference, we systematically study how network performance scales with training set size. After training, we use the neural compression $\hat{\theta}$ from fiducial and finite difference simulations to evaluate Fisher information from the neural summary, comparing it with the total information $F_{P(k)}$ of the original data. Figure 2 shows the test loss as a function of training simulation count, where the optimal performance is bounded below by
$\mathcal{L}_{CR}=\text{Tr}[\ln(\mathrm{diag}(F_{P(k)}^{-1}))].$ (6)
Figure 2: Loss ((IV.1)) on held out test data as a function of number of Quijote LH simulations. The loss asymptotes to the Cramér-Rao (C-R) bound $\mathcal{L}_{CR}$ by power law decay $\propto N^{-0.39}$ (cf. (7)). The horizontal line marks where the log loss deviates from the C-R bound by O(1). Based on our initial training set of 1500 we predict that the information extracted by the neural summary is not yet optimal and that nearly 4000 simulations are needed to be in the asymptotic regime. This prediction is verified in Fig. 5.
The figure reveals that 1500 simulations are insufficient to reach the optimal regime. While determining the precise scaling relationship from limited samples is challenging, a simple linear extrapolation in the log-log plot suggests approximately 4000 simulations are needed to saturate the information content of $P(k)$. Drawing insights from recent scaling studies of transformer models[47, 48], we hypothesize that the relationship follows a power law that asymptotes with the increasing number of simulations $N$
$\mathcal{L}=\mathcal{L}_{CR}+cN^{-\alpha},$ (7)
where $c$ is a constant and power law fitting to the test losses yields $\alpha\approx 0.39$. Although decreasing log MSE loss does not guarantee increased Fisher information across all parameters, we expect information gains for at least some parameters. In the ideal case, where loss contributions from each parameter are comparable, uncertainties should decrease uniformly across all parameters until information saturation. Figure 3 demonstrates the fraction of optimal information extracted by neural summaries at the fiducial parameter point, when trained on different numbers of simulations. More specifically, the parameter uncertainties are derived from the Fisher information matrix based on the neural summaries, calculated using (3).
Figure 3: Fisher information in neural summaries vs optimal $P(k)$. NNs trained on 500, 1250, and 1500 LH simulations show increasingly tighter parameter constraints but do not saturate the information bound of an optimal estimator based on the $P(k)$.
V.1 The Big Sobol Sequence (BSQ) Simulations
Our analysis of the Quijote Latin Hypercube simulations indicates the need for a larger training set to saturate the available cosmological information. To verify our predictions about the required simulation count and test convergence behavior, we generated a new suite of 32,768 N-body simulations. These simulations sample the parameter space ${\Omega_{m},\Omega_{b},h,n_{s},\sigma_{8}}$ using a Sobol sequence design. We utilize 16,600 of these simulations, maintaining the same 15:3:2 ratio for train, validation, and test sets as in our previous analysis. To ensure direct comparability, we employ the identical PowerSpectraNet architecture and configuration used for the Latin Hypercube simulations.
Figure 4 shows the relationship between loss and the number of simulations, demonstrating that with sufficient training data, our neural network achieves optimal summary statistics as the loss curve approaches the theoretical information bound. The BSQ simulations exhibit a slightly higher test loss compared to the LH simulations. However, the test loss follows a power law scaling of $N^{-0.36}$, remarkably consistent with our earlier prediction from the Latin Hypercube simulations, validating our methodology within 10% accuracy.
The parameter uncertainties inferred from models trained on different numbers of simulations appear in Figure 5. These results demonstrate that the information content saturates nearly 4,000 simulations, with minimal improvement beyond this point. This finding validates our earlier prediction based on power law scaling analysis of the Latin Hypercube simulations, confirming that approximately 4,000 simulations are required to extract the full information content from $P(k)$.
Figure 4: Test loss (IV.1) as a function of number of BSQ simulations. The loss follows a simple power law scaling law across 2 orders of magnitude with a power law similar to the LH simulations. The combined test loss for all parameters does not reach the optimal Fisher information computed at the fiducial point but covers the full prior range in parameter space. The loss contributions of BSQ simulations are slightly higher than the LH simulations, but Figure 5 shows that the NN achieves asymptotic optimality at the fiducial point at 4000 simulations, as predicted.
Figure 5: Fisher information in neural summaries vs optimal $P(k)$. Information starts saturating near 4000 BSQ simulations and the corresponding neural summary is nearly as informative as the optimal $P(k)$.
VI Discussion and Conclusion
This paper addresses a fundamental question in simulation-based inference for cosmology: How many simulations are required for neural networks to achieve optimal parameter inference? Our investigation of this question yields several important insights.
First, we demonstrate that current simulation suites, while impressive in scope, may be insufficient for optimal neural network training. Specifically, we show that the 2,000 Latin Hypercube simulations in the Quijote suite, while valuable, do not provide enough training data for neural networks to reach the Cramér-Rao information bound, even for the relatively simple case of the matter power spectrum $P(k)$.
Through careful analysis of neural network performance scaling, we develop a quantitative framework for predicting simulation requirements. Our initial study using Latin Hypercube simulations suggests approximately 4,000 training simulations are needed to saturate the information content from $P(k)$, based on observed convergence patterns and a simple power-law scaling assumption.
This prediction motivated the creation of our Big Sobol Sequence (BSQ), a suite of 32,768 N-body simulations that allows us to verify these scaling relationships and provide a comprehensive resource for future studies. Analysis of networks trained on the BSQ simulations confirms our earlier predictions. The test loss follows a power law decay similar to that observed in the Latin Hypercube simulations, with information content indeed beginning to saturate around 4,000 simulations. This consistency across different simulation designs suggests we have identified a fundamental aspect of neural network learning in cosmological parameter inference. Our architectural studies reveal an important relationship between network depth and training set size. While deeper networks can potentially extract more information, we find that increasing the number of training simulations often provides greater improvement than adding network layers.
These findings have significant implications for cosmological analysis. As we move toward more complex summary statistics and combine multiple probes, the simulation requirements will present a challenge save for innovations in training methodology or fast-forward models to accelerate the production of simulated training sets. Our results suggest that future simulation campaigns should carefully consider training set size requirements when allocating computational resources.
Looking forward, several questions merit further investigation. The power law scaling we observe may have deeper theoretical foundations that could inform future network design. Additionally, extending this analysis to other summary statistics and their combinations could provide valuable insights for survey analysis. The BSQ simulation suite, which we make publicly available, provides a robust foundation for such investigations. Current studies report field-level inference on weak lensing data outperforms when the NN training is conditioned on some complimentary summary i.e. power spectrum $P(k)$; this method provides optimal information when there are not enough simulations for training[49]. Although similar analysis on dark matter halo fields shows no improvement for 2000 Latin hypercube simulations. In our upcoming paper, we show a patch-based estimator to get optimal information from any survey volume without generating training simulations of present-day survey volume or requiring huge computational resources to train such a large high-resolution volume.
The methods and insights presented here extend beyond power spectrum analysis. As we prepare for next-generation surveys, understanding the requirements for optimal neural network performance becomes increasingly crucial. Our work provides both practical guidance for simulation campaigns and a framework for guiding neural information extraction in cosmological parameter inference.
VII Acknowledgements
AB acknowledges support from the Simons Foundation as part of the Simons Collaboration on Learning the Universe. The Flatiron Institute is supported by the Simons Foundation. All other post-analysis and neural training have been done on IAP’s Infinity cluster; the authors acknowledge Stéphane Rouberol for his efficient management of this facility. The BSQ simulations were generated using the rusty cluster at the Flatiron Institute. The work of FVN is supported by the Simons Foundation. BDW acknowledges support from the DIM ORIGINES 2023 INFINITY NEXT grant.
VIII Appendix
VIII.1 Effect of varying the network architecture
Neural network architecture plays a vital role in extracting information from data. While no straightforward method exists for identifying the optimal architecture, systematic exploration across different numbers of trainable parameters provides insight into the network depth and width required to obtain near-optimal results. We investigate this by examining a generic MLP architecture that varies the number of hidden layers while maintaining fixed input and output dimensions (Figure 6).
Given our task of inferring five cosmological parameters from $P(k)$, we set the input dimension to 128 and the output dimension to 5. Each hidden layer maintains the same dimension as the input layer, and we employ LeakyReLU(0.5) activation throughout. Figure 7 presents validation and test losses for MLPs with hidden layers ranging from 0 to 10, trained on varying numbers of simulations. Our analysis reveals that similar levels of Fisher information can be achieved either by increasing the training set size or by deepening the network. However, simply increasing the number of trainable parameters does not guarantee improved information extraction, since we find that deeper architectures can make optimization more challenging.
Figure 6: A generic MLP architecture of varying number of hidden layers of dimension 128. This is used to analyze the impact of the depth of the NN on the inference.
Figure 7: Loss vs depth of the MLP (i.e. number of trainable parameters). Shallower models require more realizations to reach the same performance.
Development of PowerSpectraNet
Our extensive architecture search revealed that even for the relatively simple $P(k)$ summary statistic, optimal performance requires seven fully connected layers (six hidden layers plus output) with LeakyReLU activation (Figure 1). We evaluated numerous MLPs of varying depth using 2,000 BSQ $P(k)$ samples, finding that performance plateaus at six hidden layers, with no additional information gain from deeper architectures (Figure 8). This pattern holds consistently across different training set sizes, providing two key justifications for our architectural choice. Further refinement showed improved performance when systematically varying layer widths in an expanding-contracting pattern, leading to the final PowerSpectraNet architecture.
Figure 8: Fisher information as a function of the depth of the MLP (i.e. number of trainable parameters). Information increases when the MLP is trained on enough number of simulations. Information saturates at 6 hidden layers for 2000 realizations and at this point, information does not change much even when the number of simulations is changed.
Hyperparameter Selection
Given the computational cost of hyperparameter optimization for deep architectures, we conducted an extensive hyperparameter search using a single-layer perceptron (equivalent to linear regression) on $P(k)$. This approach aligns with recent studies of large neural networks[47], which suggest hyperparameter scalability across network depths. We utilize the Weights and Biases[50] tool to optimize key hyperparameters for our neural network. Our search explores batch sizes in the range $[16,64]$, activation functions (tanh, ReLU, sigmoid, and a custom activation function (see appendix VIII.2)), optimizers (SGD, Adam, and BFGS), and learning rate schedulers (CyclicLR and StepLR). Additionally, we examine learning rates within $[0,0.01]$ and tune momentum and gamma values in the range $[0.1,1]$. This study aims to identify the optimal combination of hyperparameters to enhance model performance effectively. The optimal hyperparameter values appear in Table 1.
Hyperparameter Value
Batchsize 16
Nonlinear Activation LeakyReLU(0.5)
Optimizer Adam
Momentum 0.9
Learning Rate Scheduler StepLR
Initial Learning Rate 0.001
Stepsize 10
Gamma 0.9
Table 1: Hyperparameters are optimized for the linear model and kept fixed throughout the paper while testing different architectures.
VIII.2 Activation Function
In this section, we describe custom activation functions considered in our study other than the traditional ones to find out the best possible set of hyperparameters for neural network training. Activation functions introduce non-linearity into neural networks, allowing them to learn complex patterns. Below are the key functions used in this paper. These are piecewise continuous nonlinear functions designed to provide smooth transitions across different input ranges while mitigating issues like vanishing gradients and dead neurons. The functions are linear for $|x|>1$ and follow a cubic polynomial in the central region, ensuring smoothness and differentiability for all $x\in\mathbb{R}$.
VIII.2.1 SymmCubic
This activation function has a cubic nonlinearity and it has symmetry $\phi(-x)=-\phi(x)$ (Figure 9).
$\phi(x)=\begin{cases}x+1/3&x<-1\\ x(x^{2}+3)/6&-1\leq x\leq 1\\ x-1/3&x>1\end{cases}$ (8)
Figure 9: SymmCubic
VIII.2.2 LeakyCubic
This activation function resembles the leaky ReLU but replaces the kink at $x=0$ with a smooth cubic piece (Figure 10).
$\phi(x)=\begin{cases}x&x<-1\\ -|x|^{3}/3+x(x+2)+1/3&-1\leq x\leq 1\\ 3x&x>1\end{cases}$ (9)
Figure 10: LeakyCubic